# Neural Uncertainty Principle: A Unified View of Adversarial Fragility and LLM Hallucination

### 26 Mar 2026 v2

@@ Abstract
Adversarial vulnerability in vision and hallucination in large language models are conventionally viewed as separate problems, each addressed with modality-specific patches. This study first reveals that they share a common geometric origin: the input and its loss gradient are conjugate observables subject to an irreducible uncertainty bound. Formalizing a Neural Uncertainty Principle (NUP) under a loss-induced state, we find that in near-bound regimes, further compression must be accompanied by increased sensitivity dispersion (adversarial fragility), while weak prompt–gradient coupling leaves generation under-constrained (hallucination). Crucially, this bound is modulated by an input–gradient correlation channel, captured by a specifically designed single-backward probe. In vision, masking highly coupled components improves robustness without costly adversarial training; in language, the same prefill-stage probe detects hallucination risk before generating any answer tokens. NUP thus turns two seemingly separate failure taxonomies into a shared uncertainty-budget view and provides a principled lens for reliability analysis. Guided by this NUP theory, we propose ConjMask (masking high-contribution input components) and LogitReg (logit-side regularization) to improve robustness without adversarial training, and use the probe as a decoding-free risk signal for LLMs, enabling hallucination detection and prompt selection. NUP thus provides a unified, practical framework for diagnosing and mitigating boundary anomalies across perception and generation tasks.
@@ I Introduction
Modern neural systems have evolved far beyond their initial roles as pattern recognizers. They have become integral infrastructure in scientific discovery, industrial automation, and daily digital interaction [30, 55, 19, 64, 3, 43, 44, 26]. As these models extend across critical domains, from predicting protein structures to generating legal or medical advice, the requirement to understand their boundary behaviors shifts from a theoretical curiosity to an engineering necessity [67, 23, 22]. This fragility is most clearly manifested in two distinct yet pervasive phenomena: adversarial vulnerability in vision models, where imperceptible perturbations flip confident predictions [61, 16, 33, 38, 4, 2, 7, 6], and hallucination especially in large language models (LLMs), where generation drifts fluently from fact to fabrication [36, 35, 45, 10, 48, 29, 41, 27].
@@
Here, we take a unified view and argue that both failure modes share a geometric signature: the model enters a boundary-stress of the loss landscape. In adversarial attacks, small input perturbations can traverse locally sharp loss contours; in LLMs, weak prompt conditioning leaves a large set of continuations similarly compatible before decoding, enabling prior-driven drift. We show that these are two opposite ways of mismanaging the same “uncertainty budget”—an imbalance between input-space localization and gradient-space sensitivity.
@@
  • Neural Uncertainty Principle (NUP). We formalize a Neural Uncertainty Principle (NUP) for neural models with a well-defined scalar loss: under a boundary-emphasized (loss-induced) state, an input projection operator and its directional-derivative conjugate obey a Robertson–Schrödinger–type constraint (Theorem III.1). The constraint captures a fundamental trade-off: a model cannot be made simultaneously arbitrarily accurate on boundary-relevant samples and uniformly robust to small input perturbations.
  • From operator constraint to a computable correlation channel. A key technical contribution is that, under the loss-phase construction, the covariance term in the Robertson–Schrödinger (RS) inequality admits an exact reduction to loss-weighted statistics of the real loss gradient field (Theorem III.2, Lemma III.2). This reduction exposes an explicit input–gradient coupling channel that controls proximity to the constraint bound, and yields a practical single-backward per-sample proxy: the Conjugate Correlation Probe (CC-Probe), defined as the absolute input–gradient cosine (mean-centered for prompt embeddings).
  • Two opposite anomaly regimes under the same principle. Using the CC-Probe, we can empirically identify two opposite boundary-anomaly regimes: in vision, a persistent high-CC-Probe marks boundary-stress and concentrates adversarial fragility; in LLM prompting, anomalously low CC-Probe marks under-conditioning at prefill and correlates with elevated hallucination risk. Across settings, reliable behavior typically lies between these extremes, which we summarize as an intermediate “Goldilocks” band (Fig. 1) used as a qualitative regime indicator.
To validate the aforementioned results, we design six experiments that diagnose the two regimes and use two lightweight interventions (ConjMask, LogitReg) as mechanistic tests that manipulate the coupling channel and produce objective-specific shifts under controlled stress tests.
@@ II-A Adversarial Robustness and Accuracy–Robustness Frontier
Our work offers a complementary perspective: Proposition 1 frames this tension as a conjugate necessity arising from the NUP. As models compress features to sharpen decision boundaries, the effective feasible volume shrinks, pushing the system toward bound saturation.
@@ II-E Positioning of This Work
Existing approaches address adversarial robustness and hallucination as separate phenomena, employing modality-specific defenses (adversarial training, prompt engineering) or post-hoc diagnostics (semantic entropy, LLM-as-judge). Our work departs from this paradigm in the following manners:
@@
    Prefill signal. Unlike sampling-based uncertainty estimators, the CC-Probe requires only a single backward pass and operates before decoding (for LLMs) or attack evaluation (for vision).
@@
    Intervention design. Rather than adversarially augmenting the training set, we directly target the conjugate coupling channel that the NUP identifies as the source of fragility.
@@ III-A NUP in a Nutshell: Roadmap From Geometry to Experiments
This inequality is the only constraint required to interpret Fig. 1: the lower-left region is forbidden, i.e., one cannot simultaneously make the boundary-emphasized population extremely thin (very small $\Delta\hat{m}_{u}^{\star}$) and uniformly insensitive (very small $\Delta\hat{p}_{u}$). This is the geometric origin of adversarial vulnerability and hallucination in view of the NUP.
@@
To localize individual samples on this plane without estimating operator moments, we use a single-backward proxy for the coupling channel: the input–gradient cosine (CC-Probe), i.e., $c_{\text{img}}$ in vision and $c_{\text{prompt}}$ in LLM prefill (Eqs. (33)–(34)). It tracks the same input–gradient coupling that enters the Robertson–Schrödinger covariance term under the loss-phase construction, and empirically identifies boundary-stress in vision and under-conditioning prompts in language.
@@ III-E From operator moments to real gradients and a practical proxy
CC-Probe as a proxy for the coupling channel
@@
Motivated by Lemma III.2 and Theorem III.2, we provide the Conjugate Correlation Probe (CC-Probe, briefly) for a sample $x$ as
@@ III-F Testable predictions
The NUP motivates a simple idea: abnormal input–gradient coupling is a reliable marker of failure modes. We test this idea with the CC-Probe introduced above.
@@
As training improves clean accuracy, correctly classified images should show lower CC-Probe $c_{\text{img}}$, while misclassified/hard images remain at higher $c_{\text{img}}$ (a persistent high-cosine tail). Adversarial errors should concentrate on this high-cosine tail.
@@
Proposition 4 (Exp. 5–6 prediction: in LLM prefill, low CC-Probe means under-conditioning and higher hallucination risk).
@@ IV-B Operationalizing the Conjugate Probe
For an image input $X\in\mathbb{R}^{C\times H\times W}$, we compute the gradient of the task loss $\mathcal{L}$ with respect to the input, $P=\nabla_{X}\mathcal{L}$. Both $X$ and $P$ are flattened into 1-D vectors $x$ and $p$. Because the practical vision probe is computed on the normalized input tensor fed to the model, we use the uncentered image-space cosine; the sample-level conjugate correlation and its dataset-level average over $N$ samples are defined as:
@@ IV-C Evaluation Metrics
We treat Exp. 5 hallucination detection as a binary classification task (($1=$ hallucination, $0=$ non-hallucination)), quantified by the Area Under the ROC Curve (AUROC) with bootstrap resampling ($N_{\text{boot}}=2000$). For Exp. 6 prefill-stage prompt selection, the primary metrics are Top-1 Hit Rate (frequency with which the selected prompt belongs to the judge-best set) and Mean Regret (the average quality gap between the selected prompt and the best judge score within the same five-prompt cluster). Deterministic decoding is used for reproducibility.
@@ V-A Experimental Roadmap
Our experiments test the NUP predictions using a simple, computable proxy—the CC-Probe—across vision classification and LLM prompting. We organize Exp. 1–6 into three phases [^3] To reproduce Exp. 1–6, please refer to the released artifacts at https://doi.org/10.57760/sciencedb.29613.:
@@
V-B Exp. 1: CC-Probe Separates Correct and Incorrect Samples During Training
Figure 2: Evolution of the CC-Probe and accuracy during training. We plot evaluation accuracy (red, right axis) and the mean absolute input–gradient cosine $\bar{c}_{\text{img}}$ (left axis; Eq. (33)), computed on the held-out evaluation split every 10 epochs and reported separately for correctly classified versus misclassified samples (green vs. blue). Gradients are taken w.r.t. the standard cross-entropy loss using the ground-truth label. (a) CIFAR-10 (ResNet-18, DenseNet-121, ViT-Tiny, EfficientNet-B0). (b) Tiny-ImageNet-200 (ResNet-50, DenseNet-121, EfficientNet-B4, Swin-Tiny).
For each test input $x$ we compute the per-sample CC-Probe $c_{\text{img}}(x)$ from Eq. (33), where $p(x)=\nabla_{x}\mathcal{L}(f_{\theta}(x),y_{\text{true}})$ is the input gradient of the standard cross-entropy. We then average $c_{\text{img}}(x)$ over the subset of correctly classified samples and over the subset of misclassified samples at each epoch. Since $c_{\text{img}}$ is scale-free, the trends below are not driven by gradient magnitudes; in implementation we stabilize the cosine computation by adding a small value to the input/gradient norm denominators, rather than dropping low-gradient samples.
@@
V-C Exp. 2: Stepping Along/Against the Gradient Moves the CC-Probe as Predicted
@@
In the NUP, proximity to the constraint is governed by the effective conjugate volume $V_{c}(u)=\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}$ and the slack ratio $S_{c}(u)=V_{c}(u)/\kappa\geq 1$ (Eq. (28)). Here, the CC-Probe $\bar{c}_{\text{img}}$ is used as a proxy for the input–gradient coupling channel entering $\rho_{c}(u)$ via the loss-induced statistics; it is not itself an exact estimator of $V_{c}(u)$ or $S_{c}(u)$. Accordingly, Fig. 3 is interpreted as an empirical probe of boundary stress: +FGSM increases loss and empirically raises $\bar{c}_{\text{img}}$, which is consistent with stronger correlation pressure and higher boundary stress, whereas sufficiently small –FGSM steps reduce the measured coupling and are consistent with a relaxation toward a less stressed regime.
@@
We start from a separately trained clean CIFAR-10 checkpoint for the same architecture. For each mini-batch, we first compute a forward pass and mark samples as triggered if either the current prediction is incorrect, or the prediction confidence is below a fixed threshold (default $0.2$). For each triggered image $X$ (flattened as $x$) we compute the input gradient $P=\nabla_{X}\mathcal{L}$ (flattened as $p$). In the released implementation, this probing step intentionally switches the model to eval mode; moreover, this mode is kept for the subsequent forward passes of the same optimization step. Since the goal of ConjMask is to reduce the CC-Probe $c_{\text{img}}=|\cos(x,p)|$ (Eq. (33)), we focus on the components that contribute most to this coupling statistic.
@@
Prefill-only risk scores.
@@ Results
Overall, the results of Exp. 6 suggest that Risk-Cos can serve as a useful, decoding-free criterion for selecting among semantically equivalent prompt formulations. More importantly, this finding provides direct empirical support for the NUP account developed in this paper: within the prefill stage, stronger prompt–gradient coupling is associated with better-constrained continuations, whereas weaker coupling reflects the under-conditioning regime predicted by Proposition 4.
@@ VI Discussion
The primary theoretical value of the NUP is that it provides a fundamental boundary on what a learned system can simultaneously achieve when its behavior is examined through the joint geometry of inputs and loss sensitivities. In the Robertson–Schrödinger form (Theorem III.1), the pair $(\hat{x}_{u},\hat{p}_{u})$ obeys an irreducible constraint: the boundary-emphasized phase-plane area cannot collapse below a constant, equivalently $\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}\geq\tfrac{1}{2}$ (Lemma III.1). This plays a role analogous to a Carnot-type limit: regardless of architecture or training details, there exists a nonzero “minimum uncertainty budget” that prevents simultaneously making boundary ambiguity arbitrarily small and boundary sensitivity uniformly low. In this sense, the NUP is not merely descriptive; it states an intrinsic incompatibility that any robust-and-accurate (or faithful-and-confident) system must navigate.

### 27 Mar 2026 v3

@@ Abstract
Adversarial vulnerability in vision and hallucination in large language models are conventionally viewed as separate problems, each addressed with modality-specific patches. Here we argue that they can be interpreted through a common geometric mechanism: under a loss-induced weighting that emphasizes boundary-relevant samples, the input projection and its directional loss gradient form a conjugate pair whose dispersions obey a Robertson–Schrödinger-type lower bound—a Neural Uncertainty Principle (NUP). This bound implies that a model cannot simultaneously achieve arbitrarily sharp boundary discrimination and uniformly low sensitivity to input perturbations. Vision adversarial fragility corresponds to a near-bound regime where boundary compression inflates gradient dispersion; LLM hallucination corresponds to the opposite extreme, where weak prompt–gradient coupling leaves the continuation space under-constrained before decoding. The bound is modulated by an input–gradient correlation channel that admits exact reduction to real-gradient statistics, yielding a single-backward probe ($C$-Probe) applicable to both modalities. In vision, suppressing dominant coupling components during training improves robustness without adversarial training; in language, the same prefill-stage probe detects hallucination risk before generating any answer tokens. NUP thus reframes two isolated reliability failures as opposite sides of a single conjugate trade-off.
@@ I Introduction
Modern neural systems have evolved far beyond their initial roles as pattern recognizers. They have become integral infrastructure in scientific discovery, industrial automation, and daily digital interaction [30, 53, 19, 62, 3, 42, 43, 26]. As these models extend across critical domains, from predicting protein structures to generating legal or medical advice, the requirement to understand their boundary behaviors shifts from a theoretical curiosity to an engineering necessity [65, 21, 23]. This fragility is most clearly manifested in two distinct yet pervasive phenomena: adversarial vulnerability in vision models, where imperceptible perturbations flip confident predictions [59, 16, 34, 38, 4, 2, 7, 6], and hallucination, especially in large language models (LLMs), where generation drifts fluently from fact to fabrication [36, 33, 71, 10, 46, 29, 25, 27].
@@
Here, we take a unified view and argue that both failure modes share a geometric signature: the model enters a boundary-stressed region of the loss landscape. In adversarial attacks, small input perturbations can traverse locally sharp loss contours; in LLMs, weak prompt conditioning leaves a large set of continuations similarly compatible before decoding, enabling prior-driven drift. We show that these are two opposite ways of mismanaging the same “uncertainty budget”—an imbalance between input-space localization and gradient-space sensitivity.
@@
  • Neural Uncertainty Principle (NUP). We formalize a Neural Uncertainty Principle (NUP) for neural models with a well-defined scalar loss: under a boundary-emphasized (loss-induced) state, an input projection operator and its directional-derivative conjugate obey a Robertson–Schrödinger–type constraint (Theorem III.1). The constraint captures a principled trade-off: a model cannot be made simultaneously arbitrarily accurate on boundary-relevant samples and uniformly robust to small input perturbations.
  • From operator constraint to a computable correlation channel. A key technical contribution is that, under the loss-phase construction, the covariance term in the Robertson–Schrödinger (RS) inequality admits an exact reduction to loss-weighted statistics of the real loss gradient field (Theorem III.2, Lemma III.2). This reduction exposes an explicit input–gradient coupling channel that controls proximity to the constraint bound, and yields a practical single-backward per-sample proxy: the Conjugate Correlation Probe ($C$-Probe), defined as the absolute input–gradient cosine (mean-centered for prompt embeddings).
  • Two opposite anomaly regimes under the same principle. Using the $C$-Probe, we can empirically characterize two opposite boundary-anomaly regimes: in vision, a persistent high-$C$-Probe marks boundary-stress and concentrates adversarial fragility; in LLM prompting, anomalously low $C$-Probe marks under-conditioning at prefill and correlates with elevated hallucination risk. Across settings, reliable behavior typically lies between these extremes, which we summarize as an intermediate “Goldilocks” band (Fig. 1) used as a qualitative regime indicator.
To validate the aforementioned results, we design six experiments that diagnose the two regimes and use two lightweight interventions, ConjMask (masking high-contribution input components) and LogitReg (logit-side regularization), as mechanistic tests that manipulate the coupling channel and produce objective-specific shifts under controlled stress tests.
@@ II-A Adversarial Robustness and Accuracy–Robustness Frontier
Our work offers a complementary perspective: Proposition 1 frames this tension as a conjugate trade-off arising from the NUP. As models compress features to sharpen decision boundaries, the effective feasible volume shrinks, pushing the system toward bound saturation.
@@ II-E Positioning of This Work
Existing approaches address adversarial robustness and hallucination as separate phenomena, employing modality-specific defenses (adversarial training, prompt engineering) or post-hoc diagnostics (semantic entropy, LLM-as-judge). Our work departs from this paradigm in the following ways:
@@
    Prefill signal. Unlike sampling-based uncertainty estimators, the $C$-Probe requires only a single backward pass and operates before decoding (for LLMs) or attack evaluation (for vision).
@@
    Intervention design. Rather than adversarially augmenting the training set, we directly target the conjugate coupling channel implicated by the NUP in boundary fragility.
@@ III-A NUP in a Nutshell: Roadmap From Geometry to Experiments
This inequality is the only constraint required to interpret Fig. 1: the lower-left region is forbidden, i.e., one cannot simultaneously make the boundary-emphasized population extremely thin (very small $\Delta\hat{m}_{u}^{\star}$) and uniformly insensitive (very small $\Delta\hat{p}_{u}$). This provides the geometric intuition through which NUP relates adversarial vulnerability and hallucination.
@@
To localize individual samples on this plane without estimating operator moments, we use a single-backward proxy for the coupling channel: the input–gradient cosine ($C$-Probe), i.e., $c_{\text{img}}$ in vision and $c_{\text{prompt}}$ in LLM prefill (Eqs. (33)–(34)). It tracks the same input–gradient coupling that enters the Robertson–Schrödinger covariance term under the loss-phase construction, and empirically identifies boundary-stress in vision and under-conditioning prompts in language.
@@ III-E From operator moments to real gradients and a practical proxy
$C$-Probe as a proxy for the coupling channel
@@
Motivated by Lemma III.2 and Theorem III.2, we provide the Conjugate Correlation Probe ($C$-Probe, briefly) for a sample $x$ as
@@ III-F Testable predictions
The NUP motivates a simple idea: abnormal input–gradient coupling is a reliable marker of failure modes. We test this idea with the $C$-Probe introduced above.
@@
As training improves clean accuracy, correctly classified images should show lower $C$-Probe $c_{\text{img}}$, while misclassified/hard images remain at higher $c_{\text{img}}$ (a persistent high-cosine tail). Adversarial errors should concentrate on this high-cosine tail.
@@
Proposition 4 (Exp. 5–6 prediction: in LLM prefill, low $C$-Probe means under-conditioning and higher hallucination risk).
@@ IV-B Operationalizing the Conjugate Probe
For an image input $X\in\mathbb{R}^{C\times H\times W}$, we compute the gradient of the task loss $\mathcal{L}$ with respect to the input, $P=\nabla_{X}\mathcal{L}$. Both $X$ and $P$ are flattened into 1-D vectors $x$ and pp. Because the practical vision probe is computed on the normalized input tensor fed to the model, we use the uncentered image-space cosine; the sample-level conjugate correlation and its dataset-level average over NN samples are defined as:
@@ IV-C Evaluation Metrics
We treat Exp. 5 hallucination detection as a binary classification task ($1=$ hallucination, $0=$ non-hallucination), quantified by the Area Under the ROC Curve (AUROC) with bootstrap resampling ($N_{\text{boot}}=2000$). For Exp. 6 prefill-stage prompt selection, the primary metrics are Top-1 Hit Rate (frequency with which the selected prompt belongs to the judge-best set) and Mean Regret (the average quality gap between the selected prompt and the best judge score within the same five-prompt cluster). Deterministic decoding is used for reproducibility.
@@ V-A Experimental Roadmap
Our experiments test the NUP predictions using a simple, computable proxy—the $C$-Probe—across vision classification and LLM prompting. We organize Exp. 1–6 into three phases [^3] To reproduce Exp. 1–6, please refer to the released artifacts at https://doi.org/10.57760/sciencedb.29613.:
@@
V-B Exp. 1: $C$-Probe Separates Correct and Incorrect Samples During Training
Figure 2: Evolution of the $C$-Probe and accuracy during training. We plot evaluation accuracy (red, right axis) and the mean absolute input–gradient cosine $\bar{c}_{\text{img}}$ (left axis; Eq. (33)), computed on the held-out evaluation split every 10 epochs and reported separately for correctly classified versus misclassified samples (green vs. blue). Gradients are taken w.r.t. the standard cross-entropy loss using the ground-truth label. (a) CIFAR-10 (ResNet-18, DenseNet-121, ViT-Tiny, EfficientNet-B0). (b) Tiny-ImageNet-200 (ResNet-50, DenseNet-121, EfficientNet-B4, Swin-Tiny).
For each test input $x$ we compute the per-sample $C$-Probe $c_{\text{img}}(x)$ from Eq. (33), where $p(x)=\nabla_{x}\mathcal{L}(f_{\theta}(x),y_{\text{true}})$ is the input gradient of the standard cross-entropy. We then average $c_{\text{img}}(x)$ over the subset of correctly classified samples and over the subset of misclassified samples at each epoch. Since $c_{\text{img}}$ is scale-free, the trends below are not driven by gradient magnitudes; in implementation we stabilize the cosine computation by adding a small value to the input/gradient norm denominators, rather than dropping low-gradient samples.
@@
V-C Exp. 2: Stepping Along/Against the Gradient Moves the $C$-Probe as Predicted
@@
In the NUP, proximity to the constraint is governed by the effective conjugate volume $V_{c}(u)=\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}$ and the slack ratio $S_{c}(u)=V_{c}(u)/\kappa\geq 1$ (Eq. (28)). Here, the $C$-Probe $\bar{c}_{\text{img}}$ is used as a proxy for the input–gradient coupling channel entering $\rho_{c}(u)$ via the loss-induced statistics; it is not itself an exact estimator of $V_{c}(u)$ or $S_{c}(u)$. Accordingly, Fig. 3 is interpreted as an empirical probe of boundary stress: +FGSM increases loss and empirically raises $\bar{c}_{\text{img}}$, which is consistent with stronger correlation pressure and higher boundary stress, whereas sufficiently small –FGSM steps reduce the measured coupling and are consistent with a relaxation toward a less stressed regime.
@@
We start from a separately trained clean CIFAR-10 checkpoint for the same architecture. For each mini-batch, we first compute a forward pass and mark samples as triggered if either the current prediction is incorrect, or the prediction confidence is below a fixed threshold (default $0.2$). For each triggered image $X$ (flattened as $x$) we compute the input gradient $P=\nabla_{X}\mathcal{L}$ (flattened as $p$). In the released implementation, this probing step intentionally switches the model to eval mode; moreover, this mode is kept for the subsequent forward passes of the same optimization step. Since the goal of ConjMask is to reduce the $C$-Probe $c_{\text{img}}=|\cos(x,p)|$ (Eq. (33)), we focus on the components that contribute most to this coupling statistic.
@@
Prefill-only risk scores
@@ Results
Overall, the results of Exp. 6 suggest that Risk-Cos can serve as a useful, decoding-free criterion for selecting among semantically equivalent prompt formulations. MMore importantly, this finding provides empirical evidence consistent with the NUP account developed in this paper: within the prefill stage, stronger prompt–gradient coupling is associated with better-constrained continuations, whereas weaker coupling reflects the under-conditioning regime predicted by Proposition 4.
@@ VI Discussion
The primary theoretical value of the NUP is that it provides a principled constraint on what a learned system can simultaneously achieve when its behavior is examined through the joint geometry of inputs and loss sensitivities. In the Robertson–Schrödinger form (Theorem III.1), the pair $(\hat{x}_{u},\hat{p}_{u})$ obeys a nontrivial constraint: the boundary-emphasized phase-plane area cannot collapse below a constant, equivalently $\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}\geq\tfrac{1}{2}$ (Lemma III.1). At a high level, this resembles a limit statement: within the present operator formulation, there exists a nonzero “minimum uncertainty budget” that prevents simultaneously making boundary ambiguity arbitrarily small and boundary sensitivity uniformly low.